A solution containing 30% juice is mixed with a solution containing 10% juice to make a 100 gallons of a 12% juice mixture. How much of the juice somution was used?
Found 3 solutions by jorel555, addingup, ikleyn:
Answer by jorel555(1290)

Let n be the 30% juice. Then:
.3n+.1(100-n)=.12(100)
.3n+10-.1n=.12(100)
.2n=2
n=10 gallons of 30% juice, and 90 gallons of 10% juice needed!!!!!!!!!!!!

Answer by addingup(3677)

0.3x+0.1(100-x) = 0.12(100)
0.3x+10-0.1x = 12
0.2x = 2
x = 10 gallons of 30% solution and 90 gallons of 10% solution

Answer by ikleyn(52778)

.
A solution containing 30% juice is mixed with a solution containing 10% juice to make a 100 gallons of a 12% juice mixture.
How much of the juice solution was used?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

For detailed solutions of mixture problems see the lessons
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Typical word problems on mixtures from the archive
in this site.